# Unsupervised Cross-Domain 3D Human Pose Estimation via Pseudo-Label-Guided Global Transforms

### v1

@@ Abstract
Index Terms:
3D human pose estimation, domain adaptation, pseudo-label, global.
@@ I Introduction
Figure 1: Addressing the domain gap over camera viewpoints and locations. Rigid transformation operations (rotation and translation) are used to control pose adaptation to variations in camera viewpoints and positions. Most previous works synthesized data from source 3D poses through (a) randomly initialized rotation and translation within adversarial learning and (b) explicit computation of the translation matrices that still relied on randomly initialized rotation matrices within adversarial learning. In our method, (c) we propose explicit computation of both the rotation and translation matrices to align the source and target data.
The domain gap in 3D human pose estimation typically arises from factors such as camera viewpoint, human postures, actions, and even body sizes [16, 17, 18, 19]. To address such shortcomings, previous methods have focused mainly on data augmentation via adversarial training to increase the diversity of source data, and incorporate the augmented poses in pose estimation model training[17, 18, 20]. For example, [17] and [18] explored different ways of designing a differentiable pose augmentor that learned to change various types of factors such as the bone angle, bone length, and rigid transformation of the whole pose (rotation and translation), with [17] dealing with a single pose and [18] dealing with a pose sequence. More recently, in [20] two pose augmentors were used combined with meta-optimization to ensure the preservation of source-specific information while simultaneously exploring out-of-source distributions. However, the data augmentation process in [17, 18, 20] relied on randomly initialized GAN-based adversarial training (see Fig. 1 (a)) that only captures the source data distributions. Moreover, the discriminator in their GANs enhanced similarity between augmented and source data, so these methods were restricted to explore out-of-source distributions. Therefore, when target poses significantly deviate from the source distributions, these methods fail to address the domain gap efficiently.
@@
To effectively generate poses that simulate the target domain, it is crucial to exploit the knowledge embedded in the target dataset. However, this aspect remains heavily under-explored, with only [21] and [22] as works of significance. For example, to address the large distribution gap due to camera positions, [21] exploited the direct use of the camera’s intrinsic parameters in the computation of the transformation operation between the source and target poses. The translation matrices for human poses were explicitly calculated, while the rotations were obtained through a GAN-based augmentor (see Fig. 1 (b)) that also considered bone vector and bone length. In [22], a diffusion based model was applied to generate multiple 3D pose hypotheses, using target 2D poses for conditioning, to simulate the 3D pose distribution in the target domain. Despite their attempts, [21] and [22] do not fully leverage the latent knowledge of the target dataset and cannot ensure that the generated data remains as consistent as possible with the target dataset.
@@
In this paper, we address the domain discrepancy in 3D human pose estimation, predicated on camera viewpoints and positions, by aligning the 3D poses between the source and target camera coordinate systems through a target-domain-driven global transformation. Here, the global transformation refers to the transformation between absolute coordinates under camera coordinate systems rather than root-relative coordinates such that 2D poses can also be aligned after camera projection. Compared with augmenting rotation and translation matrices through GAN-based adversarial learning in which the parameters are randomly initialized as shown in Fig. 1(a), our target-driven global transformation is more intuitive and effective in adapting to camera viewpoints and positions (see Fig. 1(c)). The explicit computation of rotation and translation matrices make the synthesized data align with the source data in space as closely as possible.
@@
Given the 2D poses of the target dataset, we propose a Pseudo-Label Generation Module to generate 3D poses as pseudo-labels by applying constraints based on human bone length or pre-trained models [23][24]. Next, to align 3D poses between the source and target domains, we introduce a novel human body-centered coordinate system, which serves as a bridge in the rotation and translation process. Then the aligned 3D poses from the source dataset are projected into 2D using the camera intrinsic parameters of the target domain, resulting in aligned 2D poses as well. In addition to addressing variations of camera viewpoint and position, we incorporate a GAN-based Pose Augmentor to handle human posture and body size differences. The Pose Augmentor generates sequences of 3D human poses by changing bone angles and lengths. These sequences and corresponding reprojected 2D pose sequences are subsequently fed into two discriminators to refine them for better fidelity. Finally, these synthesized 2D-3D data pairs are used to fine-tune a pre-trained 3D pose estimation model. At each iteration, the pose estimator generates increasingly more accurate pseudo-labels to better adapt to the target domain.
@@
Our main contributions can be summarized as follows: (i) We propose a novel global transformation method to reduce the domain discrepancies caused by varying camera parameters. This approach is significantly resilient to camera position and viewpoint variations, without introducing any additional training parameters. (ii) We leverage the target dataset’s latent knowledge by utilizing its pseudo-labels to guide data synthesis, approximating its underlying distributions. (iii) We demonstrate the effectiveness and efficiency of the proposed method through experiments and ablations on three benchmarks, Human3.6M (H36M), MPI-INF-3DHP (3DHP) and 3DPW and achieve state-of-the-art results on H36M-3DHP and H36M-3DPW cross-dataset evaluation, outperforming the lastest state-of-the-art method by 7.2% and 11.8% respectively. Our cross-dataset evaluation results in an unsupervised setting on the target dataset even surpass the target-specific trained MixSTE model [24]. Additionally, we simulate a new viewpoint based on the 3DHP dataset to further validate the robustness of our method.
@@ II Related Work
Recent years have seen significant advances in skeleton-based 3D human pose estimation, which can be categorized into two primary approaches: (a) One-stage methods, such as [25], [26], [27], and [28], which directly predict 3D poses from input images or videos, and (b) Two-stage methods, exemplified by [29], [30], [31], and [32], which first estimate 2D poses from images or videos [33, 34, 35] and subsequently lift these to 3D. Martinez et al. [36] conducted pioneering work demonstrating that lifting ground truth 2D joint positions to 3D can be accomplished with very low errors. As 2D pose estimation techniques are quite mature [37, 38, 39], therefore here we will only focus on mainly recent methods that lift 2D poses to 3D.
PoseFormer [29] was the first purely transformer-based 3D pose estimation approach in which a spatial transformer encoded the local relationships between the 2D joints and a temporal transformer captured global dependencies across the arbitrary frames. Later, Zhang et al. [24] proposed MixSTE which had spatial and temporal transformers performing alternately to enhance feature encoding for 3D pose estimation.
More recently, diffusion models have been explored to generate multi-hypotheses for 3D pose estimation, e.g. [40][41]. These often involve a learned reverse process where noise is removed step-by-step to recover the true pose. Shan et al. [40] proposed a joint-level aggregation method based on reprojection errors from multiple hypotheses and achieved exceptional performance. Rather than directly regressing the 3D joint coordinates as in [40], the work in [41] decomposed the 3D pose into bone length and direction in the forward diffusion process to speed up gradient descent. Although these methods perform well on their training dataset, their performance tends to degrade when generalized to other datasets [22, 20].
@@ II-B Data augmentation on 3D human poses
To improve the generalisation ability of a model, previous approaches used data augmentation to increase the diversity of training data [42, 43, 44, 17, 18, 45, 46, 20]. Some methods [42] [43] focused on augmenting the original images, but more relevant to our method are those that enhanced 2D-3D pose pairs [17, 18, 45, 46, 20, 47]. Commonly in the latter methods, the data augmentor is jointly trained with the discriminator and the pose estimator end-to-end with an error-feedback training strategy. As such, the augmentor learns to augment data with guidance from the estimator and discriminator in an unsupervised way (also see Section II-C). The augmented data are used to train the pose estimator, thus improving generalization and adaption ability of the pose estimator. For example, in PoseAug [17], data augmentation was applied in terms of bone angles, lengths, and body rigid transformations. More recently, Peng et al. [20] proposed a dual-augmentor framework, which included both a weak and a strong pose augmentor. This approach preserved knowledge of source poses while exploring out-of-source distributions. [47] proposed a Biomechanical Pose Generator (BPG), a novel approach that harnessed biomechanical principles, particularly the normal range of motion, to autonomously generate diverse and realistic 3D poses. This approach did not rely on the source dataset, thus overcoming the restrictions of popularity bias.
Figure 2: Overview of the proposed method. The pseudo-3D poses from target dataset are used to guide the global transformation of the augmented 3D poses from source dataset. During the transformation, a human-centric coordinate system is proposed as a bridge. The synthesized 2D-3D pose pairs are then used to train the Discriminator $M_{d}$ and Pose Estimator $M_{e}$ jointly. The fine-tuned Pose Estimator helps generate more accurate pseudo-3D poses of the target dataset, making the framework iterative.
@@ II-C Unsupervised Domain Adaptation in 3D Pose Estimation
Unsupervised domain adaptation in 3D human pose estimation involves transferring models from the fully labeled source domain to an unlabeled target domain, with several different strategies seen in recent works. Zhang et al. [16] proposed an Inference Stage Optimization which performed geometry-aware self-supervised learning (SSL) on each single target instance and updated the 3D pose model before making predictions. Guan et al. [48] used 2D information of individual test frames and temporal constraints to fine-tune the source model on test video streams.
Recently, some works have aimed to further improve performance by synthesizing data from the source dataset in other ways [21, 22]. The synthesized data are typically designed to match the distribution of the target domain, and then used to train the pose estimator, to improve its adaption ability. For example, to address the domain gap in camera viewpoints, Chai et al. [21] exploited geometric constraints to align the root joint positions of 3D poses between the source and target domains. They also aligned the scales and root positions of 2D poses from the source and target domain using the target camera intrinsic parameters. Furthermore, they incorporated the state-of-the-art human pose estimator based on a diffusion model, i.e., D3DP [40], to generate pseudo 2D-3D pairs of the target domain [22]. Although these methods attempt to simulate the target domain, their exploration of the latent knowledge within the target domain remains insufficient or inefficient and does not guarantee the consistency between the generated data and the target dataset. This gap highlights the need for more effective methods to extract and utilize target domain knowledge to better simulate the target domain.
@@ III-A Problem Formulation
Unsupervised cross-domain 3D human pose estimation aims to adapt a pose estimation model trained on a labeled source dataset to an unlabeled target dataset. Let the human pose representation consist of $J$ skeleton joints, numbered from 00 to $J-1$ (see Fig. 3 where $J=16$). We define $\bm{P}_{2D}^{src}\in R^{2\times J}$ to denote the 2D pose and $\bm{P}_{3D}^{src}\in R^{3\times J}$ to denote the label, i.e., a 3D pose from the source dataset and, $\bm{P}_{2D}^{tgt}\in R^{2\times J}$ to represent an unlabelled 2D pose from the target dataset. Given the intrinsic parameters $\bm{K}$ of the camera from the target dataset, our method synthesizes augmented and aligned 2D and 3D pose pairs $\bm{P}^{{}^{\prime}}=(\bm{P}_{2D}^{{}^{\prime}},\bm{P}_{3D}^{{}^{\prime}})$ with

### v2

@@ Abstract
publicationid: pubid: Copyright © 2025 IEEE. Personal use of this material is permitted. However, permission to use this material for any other purposes must be obtained from the IEEE by sending an email to pubs-permissions@ieee.org.
@@ I Introduction
Figure 1: Addressing the domain gap over camera viewpoints and locations. Rigid transformation operations (rotation and translation) are used to control pose adaptation to variations in camera viewpoints and positions. Most previous works synthesized data from source 3D poses through (a) randomly initialized rotation and translation within adversarial learning and (b) explicit computation of the translation matrices but still relied on randomly initialized rotation matrices within adversarial learning. In our method, (c) we propose explicit computation of both the rotation and translation matrices to align the source and target data.
The domain gap in 3D human pose estimation typically arises from factors such as camera viewpoints, human postures, actions, and even body sizes [16, 17, 18, 19]. To address such shortcomings, some works have focussed on the task of unsupervised cross-domain 3D human pose estimation [17, 18, 19, 20]. In this setting, paired 2D and annotated 3D poses from the source domain are provided, while the target domain only offers 2D poses without corresponding 3D annotations. The objective is to train a model using all this data and then apply the model to infer 3D human poses in the target domain.
A pioneering work is [17], which proposed to augment the available training poses via adversarial training. Their augmentor learnt to change various factors, such as the bone angle, bone length, and rigid transformation of the whole pose (rotation and translation) to create augmented data. Their 3D pose estimator was better trained to generalize to other domains, given such more diverse data. In [20], two pose augmentors, combined with meta-optimization, were used to explore out-of-source distributions. This is particularly significant in cases where target poses differ significantly from source distributions. Applying a weak augmentor as an intermediary, a strong augmentor was designed to generate poses that exhibit significant deviations from the source distributions. Both [17] and [20] belong to Domain Generalization (DG) methods, in which only the source data is incorporated for the cross-domain learning.
Domain Adaptation (DA) is another type of approach for cross-domain learning by utilizing the available target domain data (i.e., input 2D poses) to improve cross-domain performance. In [18], target 2D poses were used to enforce their generated synthetic data to have similar distributions. The data augmentation process in [17, 18, 20] relied on randomly initialized GAN-based adversarial training that only captures the source data distributions (e.g. see Fig. 1 (a)). Since the synthesized data are randomly augmented, these methods are not efficient when adapting to a different domain.
@@
Some other works have considered the knowledge embedded in the target dataset to effectively generate poses that simulate the target domain. For example, to address the large distribution gap due to camera positions, [21] exploited the direct use of the camera’s intrinsic parameters to compute the transformation between the source and target poses. The translation matrices were explicitly calculated, while the rotations were obtained through a GAN-based augmentor that also considered bone vector and bone length (see Fig. 1 (b)). In [22], a diffusion-based model was applied to generate multiple 3D pose hypotheses using target 2D poses for conditioning, thus simulating the target distribution. The generated 3D pose hypotheses were then directly used as pseudo-labels to train the pose estimator. However, without proper discriminators to ensure their plausibility, the generated poses may include anatomically implausible joint configurations such as unrealistic joint angles.
@@
In this paper, we address the domain discrepancy in 3D human pose estimation, predicated on camera viewpoints and positions. Motivated by recent methods that exploit information embedded in the target dataset, we proposed to align the 3D poses between the source and target camera coordinate systems through a target-domain-driven global transformation. Here, the alignment refers to transforming the source poses through rotation and translation to minimize the average distance between the transformed source poses and the target poses over corresponding joints. In this way, the discrepancy in terms of their 3D absolute coordinates is mitigated. Additionally, the global transformation refers to the transformation between absolute coordinates under camera coordinate systems rather than root-relative coordinates such that 2D poses can also be aligned after camera projection. Compared with augmenting rotation and translation matrices through GAN-based adversarial learning in which the parameters are randomly initialized, our target-driven global transformation is more intuitive and effective in adapting to camera viewpoints and positions (see Fig. 1(c)). The explicit computation of rotation and translation matrices makes the synthesized data align with the target data in space as closely as possible.
@@
Given the 2D poses of the target dataset, we propose a Pseudo-Label Generation Module to generate 3D poses as pseudo-labels. Based on the root-relative predictions from pre-trained models [23, 24], we leverage bone length consistency and 2D reprojection constraints to infer absolute 3D joint positions as the pseudo-labels. In contrast to most prior works [17, 18, 21, 19, 20] which overlook this aspect, we explicitly explore both the generation and effective use of pseudo-labels of the target domain. In our framework, these pseudo-labels play a critical role in guiding the synthesis of data to ensure alignment with the target domain. In order to address variations in human posture and body size, we incorporate a GAN-based Pose Augmentor which generates sequences of 3D human poses by changing bone angles and lengths. Next, to align 3D poses between the augmented source domain and the target domain, we introduce a novel human body-centered coordinate system, which serves as a bridge in the rotation and translation process. Then the aligned 3D poses are projected into 2D using the camera intrinsic parameters of the target domain, resulting in aligned 2D poses as well. These sequences and corresponding re-projected 2D pose sequences are subsequently fed into two discriminators for better fidelity. These synthesized 2D-3D data pairs are also used to fine-tune a pre-trained 3D pose estimation model. At each iteration, the pose estimator generates increasingly more accurate pseudo-labels to better adapt to the target domain. The above modules are integrated into a unified framework, cohesively bridging the domain gap and substantially improving 3D pose estimation performance when adapted to the target domain.
@@
Our main contributions can be summarized as follows: (i) We propose a novel global transformation method to reduce the domain discrepancies caused by varying camera parameters. This approach is significantly resilient to camera position and viewpoint variations, without introducing any additional training parameters. (ii) We exploit the target dataset’s latent knowledge and generate 3D pseudo-labels that better approximate the target distribution to guide the global transformation. (iii) We demonstrate the effectiveness and efficiency of the proposed method through experiments and ablations on three benchmarks, Human3.6M (H3.6M), MPI-INF-3DHP (3DHP), 3DPW and achieve state-of-the-art results on H3.6M-3DHP and H3.6M-3DPW cross-dataset evaluation, outperforming the lastest state-of-the-art method by 7.2% and 11.8% respectively. Our cross-dataset evaluation results in an unsupervised setting on the target dataset even surpass the target-specific trained MixSTE model [24]. Additionally, we simulate a new viewpoint based on the 3DHP dataset to further validate the robustness of our method.
@@ II Related Work
Recent years have seen significant advances in skeleton-based 3D human pose estimation, which can be categorized into two primary approaches: (a) One-stage methods, such as [25], [26], [27], and [28], which directly predict 3D poses from input images or videos; (b) Two-stage methods, exemplified by [24, 29, 30, 31, 32], which first estimate 2D poses from images or videos [33, 34, 35] and subsequently lift these to 3D. Martinez et al. [36] conducted a pioneering work demonstrating that lifting ground truth 2D joint positions to 3D can be accomplished with very low errors. As 2D pose estimation techniques are quite mature [37, 38, 39], we only focus on recent methods that lift 2D poses to 3D.
In [30], a novel Graph Convolutional Network was proposed for 3D HPE. Weight modulation and affinity modulation were designed to assign unique modulation vectors to nodes and update the graph structure respectively, thereby improving performance without increasing model size. With the rise of transformers, PoseFormer [29] was the first purely transformer-based 3D pose estimation approach in which a spatial transformer encoded the local relationships between the 2D joints and a temporal transformer captured global dependencies across the arbitrary frames. Later, Zhang et al. [24] proposed MixSTE, which took a sequence of 2D keypoints as input and predicted 3D poses. It alternately applied spatial and temporal transformer blocks to capture spatio-temporal dependencies. Li et al. [31] introduced MHFormer, which generated multiple pose hypotheses via a multi-hypothesis transformer to better model temporal uncertainty. Tang et al. [32] further proposed a two-pathway block that modelled spatial and temporal information in parallel.
More recently, diffusion models have been explored to generate multiple hypotheses for 3D pose estimation, e.g. [40][41]. These often involve a learned reverse process where noise is removed step-by-step to recover the true pose. Shan et al. [40] proposed a joint-level aggregation method based on reprojection errors from multiple hypotheses and achieved exceptional performance. Rather than directly regressing the 3D joint coordinates as in [40], the work in [41] decomposed the 3D pose into bone length and direction in the forward diffusion process to speed up gradient descent. Although these methods perform well on their training dataset, their performance tends to degrade when generalized to other datasets [22, 20].
@@ II-B Data augmentation on 3D human poses
To improve the generalization ability of a model, previous approaches used data augmentation to increase the diversity of training data [42, 43, 44, 45, 17, 18, 46, 47, 20]. Some methods [42, 43, 45] focused on augmenting the original images, but more relevant to our method are those that enhanced 2D-3D pose pairs [17, 18, 46, 47, 20, 48]. In the latter methods, the data augmentor is jointly trained with the discriminator and the pose estimator end-to-end with an error-feedback training strategy. As such, the augmentor learns to augment data with guidance from the estimator and discriminator in an unsupervised way (also see Section II-C). The augmented data are used to train the pose estimator, thus improving the generalization and adaptation ability of the pose estimator. For example, in PoseAug [17], data augmentation was applied in terms of bone angles, lengths, and body rigid transformations. More recently, Peng et al. [20] proposed a dual-augmentor framework, which included both a weak and a strong pose augmentor. This approach preserved knowledge of source poses while exploring out-of-source distributions. [48] proposed a Biomechanical Pose Generator (BPG), a novel approach that harnessed biomechanical principles, particularly the normal range of motion, to autonomously generate diverse and realistic 3D poses. This approach did not rely on the source dataset, thus overcoming the restrictions of popularity bias.
Figure 2: Overview of the proposed method. The pseudo-3D poses from target dataset are used to guide the global transformation of the augmented 3D poses from source dataset. During the transformation, a human-centric coordinate system is proposed as a bridge. The synthesized 2D-3D pose pairs are then used to train the discriminators $M_{d}^{2D}$, $M_{d}^{3D}$ and Pose Estimator $M_{e}$ jointly. The fine-tuned Pose Estimator helps generate more accurate pseudo-3D poses of the target dataset, making the framework iterative. In the Global Transformation module $M_{t}$, the symbols $C_{s}$ and $C_{t}$ represent the camera coordinate systems for the source ($s$) and target ($t$) domains, respectively. $H$ denotes the human-centric coordinate system. $R_{1}$, $R_{2}$ represent the rotation matrices among coordinate systems and $T_{1}$, $T_{2}$ represent the translation matrices.
@@ II-C Unsupervised Domain Adaptation in 3D Pose Estimation
Unsupervised domain adaptation in 3D human pose estimation involves transferring models from the fully labelled source domain to an unlabelled target domain, with several different strategies seen in recent works. Zhang et al. [16] proposed an Inference Stage Optimization which performed geometry-aware self-supervised learning (SSL) on each single target instance and updated the 3D pose model before making predictions. Guan et al. [49] used 2D information of individual test frames and temporal constraints to fine-tune the source model on test video streams.
Recently, some works have aimed to further improve performance by synthesizing data from the source dataset in other ways [21, 22]. The synthesized data are typically designed to match the distribution of the target domain, and then used to train the pose estimator, to improve its adaptation ability. For example, to address the domain gap in camera viewpoints, Chai et al. [21] exploited geometric constraints to align the root joint positions of 3D poses between the source and target domains. While the alignment of the root joint was achieved through the explicit computation of the translation matrix, the alignment of the remaining joints relied on the rotation matrix, which was still learnt via GAN-based methods. In [22], the diffusion-based human pose estimator D3DP [40] was deployed to generate pseudo 2D-3D pairs of the target domain, which were directly used to train the pose estimator. However, without checking for plausibility, their generated 3D pose hypotheses may exhibit unreasonable joint angles. The above methods attempted to simulate the target domain via exploration of the latent knowledge within the target domain and achieved some improvements. Inspired by them, we aim to more fully exploit target domain information by generating pseudo-labels to guide synthetic data while ensuring their plausibility.
@@ III-A Problem Formulation
Unsupervised cross-domain 3D human pose estimation aims to adapt a pose estimation model trained on a labelled source dataset to an unlabelled target dataset. Let the human pose representation consist of $J$ skeleton joints, numbered from 0 to J1J-1. We define $\bm{P}_{2D}^{src}\in R^{2\times J}$ to denote the 2D pose and $\bm{P}_{3D}^{src}\in R^{3\times J}$ to denote the label, i.e., a 3D pose from the source dataset and, $\bm{P}_{2D}^{tgt}\in R^{2\times J}$ to represent a 2D pose from the target dataset which does not have any 3D pose label. Given the intrinsic parameters $\bm{K}$ of the camera from the target dataset, our method synthesizes augmented and aligned 2D and 3D pose pairs $\bm{P}^{{}^{\prime}}=(\bm{P}_{2D}^{{}^{\prime}},\bm{P}_{3D}^{{}^{\prime}})$ with
@@
III-B Overview of the Proposed Method
Our proposed framework aims to synthesize new 2D-3D pose pairs from the labelled source domain and the unlabelled target domain, and then use them to train a pose estimator to predict the 3D human pose in the target domain. Among the various factors contributing to the domain gap, camera viewpoints and positions play a significant role by affecting the global positions of the poses. Thus, the key challenge to be met is how to make the augmented data adequately match the distribution of the target dataset. To achieve this, we explore the latent knowledge of the target dataset as follows (also illustrated in Fig. 2). A pseudo-label generation module $M_{p}$ is proposed to generate pseudo-3D poses for the target dataset based on the Pose Estimator $M_{e}$, which is pre-trained on the source domain. For the source dataset, we employ an augmentation module, Pose Augmentor $M_{a}$, that creates more diverse samples to better account for motions and body sizes. The 2D and 3D poses from both domains are then spatially aligned in a Global Transformation module $M_{t}$ using a human-centric coordinate system as a bridge. These aligned 2D and 3D poses are then used to train the Pose Discriminators $M_{d}^{2D}$, $M_{d}^{3D}$, and the Pose Estimator $M_{e}$ jointly, while the discriminators are trained adversarially to maintain both the diversity and plausibility of synthesized data. Finally, the fine-tuned Pose Estimator $M_{e}$ is applied to $M_{p}$ to generate updated 3D pseudo-labels of the target dataset again. Since the Global Transformation module is guided by pseudo-labels, updated pseudo-labels enhance alignment between synthesized and target data. Therefore, it may be possible to boost the pose estimation results by iterating the above process. We explore this in our ablations.
III-C Pseudo-Label Generation
For the target dataset, we generate 3D pseudo-labels by inferring absolute 3D coordinates of human skeleton joints in the camera coordinate system from the available 2D poses. These 3D pseudo-labels can be obtained using a pre-trained pose estimation method given 2D inputs.
Given a 2D pose $\bm{P}_{2D}^{tgt}=[\bm{x}^{tgt},\bm{y}^{tgt}]^{T}$ from the target dataset, we obtain the root-relative 3D coordinates $\hat{\bm{P}}_{3D}^{tgt}=[\hat{\bm{X}}^{tgt},\hat{\bm{Y}}^{tgt},\hat{\bm{Z}}^{tgt}]^{T}$ using the pre-trained pose estimator $M_{e}$ on the source dataset. The predicted 3D coordinates of the joints $(\hat{X_{j}}^{tgt},\hat{Y_{j}}^{tgt},\hat{Z_{j}}^{tgt}),j=\left\{0,...,J-1\right\}$ are root-relative 3D coordinates with the root $(\hat{X_{0}}^{tgt},\hat{Y_{0}}^{tgt},\hat{Z_{0}}^{tgt})$ set to $(0,0,0)$. To obtain the absolute coordinates of the root, we apply a two-stage optimization process. In the first stage, given the 2D pose and camera intrinsic parameters $\bm{K}$, the 3D pose can be inferred by assuming an initial depth value.
The corresponding bone lengths are then derived from the estimated 3D pose. To ensure anatomical consistency, we minimize the discrepancy between the estimated bone lengths and their averaged values using least squares, and thus, we obtain the 3D absolute coordinates of all joints $\tilde{\bm{P}}_{j}=({\tilde{X}_{j}}^{tgt},{\tilde{Y}_{j}}^{tgt},{\tilde{Z}_{j}}^{tgt}),j=\left\{0,...,J-1\right\}$. In the second stage, we minimize the error between reprojected 2D pose from $\hat{\bm{P}}_{3D}^{tgt}+\bm{P}_{r}$ and real 2D pose $\bm{P}_{2D}^{tgt}$ using the least squares method again, where $\bm{P}_{r}$ is the absolute coordinates of the root joint and its initial value is set to root joint $\tilde{\bm{P}}_{0}$. Finally, we arrive at the estimated absolute 3D coordinates of all joints $\tilde{\bm{P}}_{3D}^{tgt}$ as the 3D pseudo-labels of the target dataset.
There are no training parameters in $M_{p}$, while the absolute 3D coordinates are obtained by applying bone-length and 2D projection constraints through the least squares optimization process. Based on the above pseudo-label generation method, we generate 3D pseudo-labels that approximate the target distribution, exploiting the latent knowledge of the target dataset. These pseudo-labels play a pivotal role in guiding the subsequent global transformation process.
III-D Pose Augmentor
The poses in the source dataset are augmented using the bone augmentation process from AdaptPose [18]. This enables our Pose Augmentor $M_{a}$ to modify bone angles across frames, introducing greater motion diversity. As a result, the augmented poses capture a wider range of actions. Additionally, bone lengths are adjusted to accommodate variations in human body proportions, enhancing the model’s adaptability to different body sizes.
Given a sequence of 3D poses $\left\{\bm{P}_{3D}^{src}\right\}_{n=0}^{N-1}$ from the source dataset, our augmentor converts them into their corresponding bone vectors $\left\{\bm{B}^{src}\right\}_{n=0}^{N-1}\in R^{N\times(J-1)\times 3}$. Each original bone vector $b^{l}_{n}\in R^{1\times 3}$ (i.e., the $l$-th bone vector from $\bm{B}^{src}$ in $n$-th frame) is rotated about an axis $\bm{w}^{l}\in R^{1\times 3}$ by a certain rotation angle $\theta^{l}$ proportional to the number of frames. This guarantees that the bone changes are temporally plausible and consistent. Additionally, to adjust the bone length, the pose augmentor scales it using a ratio parameter $\lambda^{l}$, which systematically modifies the bone length while preserving the structural coherence of the motion.
To predict the parameters ($\bm{w}\in R^{(J-1)\times 3},\theta\in R^{(J-1)\times 1},\lambda\in R^{(J-1)\times 1}$), following [18], the input is formed by concatenating a Gaussian-noise vector with the 3D pose of the middle frame as the representative of the whole sequence. The Gaussian noise introduces randomness and diversity for feature extraction. Subsequently, four fully connected layers followed by batch normalization and a leaky ReLU are used to extract features and learn these parameters [18].